OpenStudy (anonymous):

4(3-x)^4/3-5=59 please help

OpenStudy (andijo76):

x=−5

OpenStudy (anonymous):

can you show me the steps please

OpenStudy (anonymous):

when you get to 3-x=8 what do you do ?

OpenStudy (jdoe0001):

\(\bf \cfrac{4(3-x)^4}{3}-5=59\implies \cfrac{4(3-x)^4}{3}=59+5\implies \cfrac{4(3-x)^4}{3}=64\\ \quad \\ 4(3-x)^4=64\cdot 3\implies 4(3-x)^4=192\implies (3-x)^4=\cfrac{192}{4}\\ \quad \\ \textit{then take }{\large \sqrt[4]{\qquad }}\textit{to both sides and 3 and multiply by -1}\)

OpenStudy (jdoe0001):

\(\textit{then take }{\large \sqrt[4]{\qquad }}\textit{to both sides and subtract 3 and multiply by -1}\)

OpenStudy (anonymous):

thank you!